OpenStudy (anonymous):

67 x 56 x formula 4 x 5

OpenStudy (anonymous):

Simply use the quadratic formula\[(-b \pm \sqrt{b^2-4ac})/2a\] Where your equation becomes: \[ax^2+bx+c=0\] Don't forget to plug in the plus and minus of the quadratic formula.

OpenStudy (anonymous):

In this example it seems the equation doesn't have any roots. Hence you do not get any values from the quadratic equation. Are you sure the equation is indeed : 3x^2 - x + 6 = 0